skipping your daily Starbucks fix would save you $180 per month. If you invested those savings in a retirement account earning 0.25% per month compound interest, you would have an additional $166,691 saved up by time you retire. Assuming you will live 20 years after retiring, how much additional monthly income will this provide you in your retirement
Answer #1

You have an additional total sum of $166,691 saved in your retirement account.

Since there are 12 months in an year, the number of months in 20 years = 12 × 20 = 240 months

Additional monthly retirement income = 166691 / 240 = $ 694.54 per month